# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v3

@@ Abstract.
With over 760 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models. Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated. In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer. OODTE adopts a straightforward yet powerful differential testing, fault localization, and evaluation methodology, which can be readily adapted for use with other compiler optimizers. Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization. When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity. We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis. Our evaluation revealed that 9.29.2% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies. Additionally, 3030% of classification models and 16.616.6% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization. In total, OODTE uncovered 15 issues, 14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly. All findings were reported to the ONNX Optimizer development team. OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
@@ 1. Introduction
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (speedisallyouneed, ; CompilerGym, ). Widely-used AI compilers and frameworks, such as Apache TVM (tvm, ) and MLIR (lattner2020mlircompilerinfrastructureend, ), incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
Models can also be represented using industry-standard formats, such as ONNX (onnxsite, ). By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer (onnxoptimizer, ) serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a suite that detects and localizes faults in the ONNX Optimizer.
@@
OODTE solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations. In particular, OODTE employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
@@
To evaluate the effectiveness of OODTE, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation. All models were obtained from the official ONNX Model Hub (onnxmodelhub, ). Each model was optimized using the ONNX Optimizer and assessed on standard benchmark datasets to compare accuracy before and after optimization. When discrepancies were detected, we applied pass-by-pass optimization to isolate the specific transformations responsible for the faults.
We focus on the ONNX Optimizer due to its impact and popularity (over 760 GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (onnxsimplifier, ), which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. With OODTE, we aim to enhance the effectiveness of fault localization and detection in the optimizer. However approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
By conducting experiments on our case studies, we observed output discrepancies between the original and optimized models in $34$ of $130$ cases (26.126.1%). We also encountered issues with $12$ model instances (9.29.2%) using the primary optimizer passes (fuse and eliminate) where the optimizer crashed or produced invalid models. Through analysis, we identified 15 distinct bugs, only one of which had been partially reported before (without a full explanation or reproduction steps). For this case, we provided additional insights. All bugs were reported to the official ONNX Optimizer repository with detailed descriptions for further resolution. OODTE effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer. These manifested as (1) optimizer crashes, (2) malformed models, and (3) valid models with output label inconsistencies compared to their unoptimized counterparts.
@@ 2.1.1. Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (softmax, ).
@@ 2.1.3. Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers (attention, ), use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
@@ 2.2. Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (tvm, ; lattner2020mlircompilerinfrastructureend, ) incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
@@ 2.3. The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify the graph structure to ensure compatibility; for example, rewrite_input_dtype adjusts input data types. In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_pre-
dict
, which isolate the initializer and inference subgraphs, respectively. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
@@ 2.4. Model Output Comparison Methods
Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts. In particular:
@@
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (kendall, ), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection & Semantic Segmentation models::
Ea-
ch prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of $0.5$, $0.75$, and $0.9$, with the original model as the reference.
@@
We convert token outputs to plain text and use the BLEU Score (BLEU, ) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
@@ 3. System Architecture & Methodology
Figure 1. Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE and its modules is shown in Figure 1. Its workflow is straightforward: load models, optimize them, execute both original and optimized versions across the dataset, and compare the results. If output discrepancies are detected, OODTE iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference. All results are stored in report files for analysis and fault localization.
OODTE consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. It loads model configuration settings either (1) from a local configuration file or (2) by fetching them from the ONNX repository from which the models are sourced, both managed by the Model Orchestrator Module. The modules in Figure 1 are detailed in the following sections.
@@ 3.1. Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading models from an online repository. At the moment, OODTE supports the official ONNX Model Hub (onnxmodelhub, ), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
@@ 3.2. Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model, applied in a user-defined order specified in the system configuration. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This feature is essential throughout the OODTE pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth. To identify the pass responsible for problematic behavior, OODTE first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified. Once a model is optimized with either setting, it then gets propagated to the Runner model for inference.
@@ 3.3. Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime (onnxruntime, ), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (datasets, ) package.
@@ 3.4. Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (kendall, ). For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (IoUBoundingBox, )), Precision, Recall, and F1 Score (F1Score, ) given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU (BLEU, ) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.

### v5

# DiTOX: Fault Detection and Localization in the ONNX Optimizer

@@ Abstract.
The ONNX Optimizer, part of the official ONNX repository and widely adopted for graph-level model optimizations, is used by default to optimize ONNX models. Despite its popularity, its ability to preserve model correctness has not been systematically evaluated. We present DiTOX, an automated framework for comprehensively assessing the correctness of the ONNX Optimizer using differential testing, fault localization, and evaluation techniques that generalize to other compiler optimizers. DiTOX applies optimization passes to a corpus of ONNX models, executes both original and optimized versions on user-defined inputs, and detects discrepancies in behavior or optimizer failures. When divergences are observed, DiTOX isolates the responsible optimization pass through iterative, fine-grained analysis. We evaluated DiTOX on 130 models from the ONNX Model Hub spanning vision and language tasks. We found that 9.2% of model instances crashed the optimizer or produced invalid models under default settings. Moreover, output discrepancies occurred in 30% of classification models and 16.6% of object detection and segmentation models, while text-based models were largely robust. Overall, DiTOX uncovered 15 issues—14 previously unknown—affecting 9 of the 47 optimization passes as well as the optimizer infrastructure. All issues were reported to the ONNX Optimizer developers. Our results demonstrate that DiTOX provides a simple and effective approach for validating AI model optimizers and is readily extensible beyond ONNX.
Software Engineering, Software Testing, Fault Localization, Differential Testing, ONNX Optimizer
conference: Proceedings of the 35th ACM SIGPLAN International Conference on Compiler Construction; January 31 – February 1, 2026; Sydney, NSW, Australiabooktitle: Proceedings of the 35th ACM SIGPLAN International Conference on Compiler Construction ($C$ ’26), January 31 – February 1, 2026, Sydney, NSW, Australiaccs: Software and its engineering Empirical software validationccs: Software and its engineering Software defect analysisccs: Software and its engineering Operational analysis
@@ 1. Introduction
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (Chen et al., 2023; Cummins et al., 2022). Widely-used AI compilers and frameworks, such as Apache TVM (Chen et al., 2018) and MLIR (Lattner et al., 2020a), incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy. Models can also be represented using industry-standard formats, such as ONNX (43). By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer (40) serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. To this end, we propose the Differential Testing of the ONNX Optimizer (DiTOX) [^1] The source code of DiTOX, along with usage instructions, is available at https://github.com/luludak/DiTOX. (Louloudakis, 2025), a suite that detects and localizes faults in the ONNX Optimizer.
@@
DiTOX solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations. In particular, DiTOX employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
@@
To evaluate the effectiveness of DiTOX, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation. All models were obtained from the official ONNX Model Hub (39). Each model was optimized using the ONNX Optimizer and assessed on standard benchmark datasets to compare accuracy before and after optimization. When discrepancies were detected, we applied pass-by-pass optimization to isolate the specific transformations responsible for the faults.
We focus on the ONNX Optimizer due to its impact and popularity (over 770 GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (42), which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. With DiTOX, we aim to enhance the effectiveness of fault localization and detection in the optimizer. However approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
By conducting experiments on our case studies, we observed output discrepancies between the original and optimized models in $34$ of $130$ cases (26.126.1%). We also encountered issues with $12$ model instances (9.29.2%) using the primary optimizer passes (fuse and eliminate) where the optimizer crashed or produced invalid models. Through analysis, we identified 15 distinct bugs, only one of which had been partially reported before (without a full explanation or reproduction steps). For this case, we provided additional insights. All bugs were reported to the official ONNX Optimizer repository with detailed descriptions for further resolution. DiTOX effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer. These manifested as (1) optimizer crashes, (2) malformed models, and (3) valid models with output label inconsistencies compared to their unoptimized counterparts.
@@ 2.1.1. Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (Pearce et al., 2021).
@@ 2.1.3. Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers (Vaswani et al., 2017), use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
@@ 2.2. Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (Chen et al., 2018; Lattner et al., 2020a) incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
@@ 2.3. The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes [^2] The complete list of passes, each accompanied by a brief description, can be found at https://github.com/luludak/DiTOX/blob/master/passes.txt. grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify the graph structure to ensure compatibility; for example, rewrite_input_dtype adjusts input data types. In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_predict, which isolate the initializer and inference subgraphs, respectively. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
@@ 2.4. Model Output Comparison Methods
Based on the model and output type, DiTOX uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts. In particular:
@@
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (Kendall, 1938), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection and Semantic Segmentation models::
Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of $0.5$, $0.75$, and $0.9$, with the original model as the reference.
@@
We convert token outputs to plain text and use the BLEU Score (Papineni et al., 2002) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
@@ 3. System Architecture & Methodology
Figure 1. Architecture of DiTOX with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of DiTOX and its modules is shown in Figure 1. Its workflow is straightforward: load models, optimize them, execute both original and optimized versions across the dataset, and compare the results. If output discrepancies are detected, DiTOX iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference. All results are stored in report files for analysis and fault localization.
DiTOX consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. It loads model configuration settings either (1) from a local configuration file or (2) by fetching them from the ONNX repository from which the models are sourced, both managed by the Model Orchestrator Module. The modules in Figure 1 are detailed in the following sections.
DiTOX currently isolates bugs at the level of individual passes, relying on compiler error messages to identify the cause. For more effective fault localization, future enhancements can explore data flow analysis and delta debugging within the identified pass to identify the root cause.
@@ 3.1. Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading models from an online repository. At the moment, DiTOX supports the official ONNX Model Hub (39), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
@@ 3.2. Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model, applied in a user-defined order specified in the system configuration. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). DiTOX supports the utilization of these passes for testing, in order to explore all optimizer operations. This feature is essential throughout the DiTOX pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth. To identify the pass responsible for problematic behavior, DiTOX first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified.
Once a model is optimized with either setting, it then gets propagated to the Runner model for inference.
@@ 3.3. Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime (41), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. DiTOX supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (8) package.
@@ 3.4. Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, DiTOX supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (Kendall, 1938). For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (Števuliáková and Hurtik, 2023)), Precision, Recall, and F1 Score (Christen et al., 2023) given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU (Papineni et al., 2002) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition DiTOX is build in a manner that new comparators can be added, in cases that a new category of models is tested.